OpenStudy (anonymous):

Explain how you would find the exact value of 3 --- √8

OpenStudy (zepp):

Multiply by sqrt of 8 to both numerator and denominator \(\large \frac{3\sqrt{8}}{8}\) Rewrite that as \(\large \frac{3}{8}\sqrt{8}\) Since a square root gives two answers \(\large \pm\sqrt{8}*\frac{3}{8}\)